# Integration of TinyML and LargeML: A Survey of 6G and Beyond

### v1

@@ VII Conclusion
TinyML–LargeML integration is set to revolutionize the development of advanced AI solutions in 6G and future networks. These technologies address the critical issues of resource allocation, network management, security, and privacy. This survey provides a comprehensive examination of the convergence of TinyML and LargeML technologies, beginning with a detailed review of design considerations, operational mechanisms, and performance evaluation methods. The survey then investigates the motivations and technical requirements driving the convergence of these systems toward 6G and beyond, presenting efficient bidirectional integration strategies and highlighting application scenarios. This extensive review identifies the challenges persistent throughout the literature and outlines numerous promising research directions, including standardization, resource orchestration, advanced security and privacy-preserving techniques, real-time and lightweight AI processing, and the adoption of AI-native and collective AI paradigms. Sustained research in this domain is essential. This survey aims to support and inform the ongoing work in providing a foundation for future innovations and advancements in 6G systems.

### v2

@@ VII Concluding Remarks
TinyML–LargeML integration is set to reshape AI design for 6G by enabling efficient, intelligent solutions to challenges in resource allocation, network management, security, and privacy. This survey reviews the foundations of this convergence, covering design principles, operational mechanisms, and performance evaluation while outlining the motivations and technical requirements driving bidirectional integration and its emerging application scenarios across 6G ecosystems. This survey also identifies key open challenges and promising research directions, including O‑RAN standardization, cross‑layer resource orchestration, advanced security and privacy‑preserving methods, real‑time lightweight AI, AI-enabled 6G communication-sensing-computing, the shift toward AI‑native and collective AI architectures, and application prospects and outlooks for 6G and AI. Continued progress in these areas is essential, and the insights provided here aim to guide and support future 6G‑enabled intelligent systems.